# Comprehensive Evaluation on Lexical Normalization: Boundary-Aware Approaches for Unsegmented Languages

### v1

@@ D.3 Comparison of Generative Approaches
D.4 Results for Each Category
Table 15 shows the JMLN test performance (recall) of the three models evaluated in §7.5 for each domain. Consistent with the domain-specific evaluation results in §7.5, Sarashina achieved the highest recall across all categories. All three models showed the same recall order across the four categories, and the consistently lowest recall of the fourth category, i.e., typos, once again highlights the difficulty of correcting them.
Category # DeBERTa T5 Sara2.2
@@ D.4 Results for Each Category
Table 15: JMLN test recall of representative models (DeBERTa-large, T5-large, and Sarashina2.2-3b) for each category.
D.5 Prediction Examples
@@
Table 16: Detailed results of encoder-only models on the JMLN dev sets. P-Seg, F-Seg, and F-Seg-POS indicate the Part-Seg, Full-Seg, and Full-Seg-POS approaches.
@@
Table 17: Detailed results of encoder-decoder and decoder-only models on the JMLN dev sets.
[…]
@@
[…]
      (The high is like 25{CJK}UTF8ipxm℃ or something, right?)
[…]
@@
[…]

### v2

@@ E.4 Comparison of Generative Approaches
Notably, we also trained a character-level Transformer based on the T5 architecture (the same size as T5-efficient-tiny [^25] https://huggingface.co/google/t5-efficient-tiny) from scratch with full JMLN training set, but it failed to reach adequate performance ($\approx$14 chrF on the dev set).
E.5 Informal Span Detection Accuracy for Thai
Table 18 shows the performance of the informal word span detection task for selected Thai models on the VISTEC test set. Consistent with the Japanese results in §7.6, the encoder-only model exhibited larger precision/recall differences between the two tasks, while the recall differences remained similar across the three model types. Overall, the high-precision encoder-only model and the high-recall encoder-decoder and decoder-only models yielded in F0.5 scores at a comparable level.
E.6 Results Across Category
Table 19 shows the JMLN test performance (recall) of the three models evaluated in §7.5 for each variant-form type category. Consistent with the domain-specific evaluation results in §7.5, Sarashina2.2-3B achieved the highest recall across all categories. All three models showed the same recall order across the four categories, and the consistently lowest recall of the fourth category, i.e., typos, once again highlights the difficulty of correcting them.
Category # DeBERTa T5 Sarashina
@@ E.6 Results Across Category
Table 19: JMLN test recall of representative models (DeBERTa-large, T5-large, and Sarashina2.2-3B) for each variant-form type category.
E.7 Prediction Examples
@@
Table 20: Results of encoder-only models with different approaches (P-Seg: Part-Seg, F-Seg: Full-Seg, and F-Seg-POS: Full-Seg-POS) on the JMLN dev sets. For each model, the best score among its variants is underlined.
@@
Table 21: Normalization results of encoder-decoder and decoder-only models on the JMLN dev sets. For each model, the best score among its variants is underlined.
(a) RK-ICB: $(\cdots)$ 効果はわからない〜〜Pos:Unk
@@
(f) NU: $(\cdots)$ 最高気温25度とかっしょうPos:Unk?
      (The high is like 25℃ or something, right?)
Gold $\rightarrow{でしょう}$ (right?)
@@
Table 22: Example original text (fragment) and corresponding model outputs. For each row of gold-standard and model output, columns 2–5 indicate: (1) original $\rightarrownormalized string$, (2) gloss of the normalized string, (3) error type, (4) manual validity judgement by the first author ($\checkmark$: valid, $\triangle$: questionable, ✗: invalid). Non-standard words in the original text are underlined; if positive instances, they are marked with a subscript $\mathsf{Pos:Unk}$, and if negative, with $\mathsf{Neg:Unk}$ (Unk indicates that the word did not appear in the training set). Predicted normalized strings are marked with subscripts $\mathsf{InL}$ or $\mathsf{OOL}$, indicating whter the form is included or not included in the UniDic lexicon.